Circular No. 5782

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1993L IN IC 5270 M. Della Valle and E. Cappellaro, European Southern Observatory,

report: “Analysis of a spectrogram (range 380-800 nm, resolution ∼ 1 nm), obtained with the New Technology Telescope at La Silla on May 1.4 UT, shows this to be a type-Ia supernova. The spectrum is dominated by absorption lines of Si II (635.5, 597.2, 564.0, 545.4, 412.9 nm). A preliminary measurement of the expansion velocity deduced from the minimum of the Si II absorptions yields ∼ 16 000 ± 500 km/s. Additional absorption of Mg II (448.1 nm) is blueshifted by 13 000 km/s. The spectrum closely resembles that of SN 1989B (Barbon et al. 1990, A.Ap. 237, 79) around maximum.”

N. Suntzeff, Cerro Tololo Interamerican Observatory, communicates: “CCD spectrophotometry taken with the CTIO 1.5-m telescope by V. Robledo-Rella on May 2.40 UT confirms that the bright object reported by Evans is a supernova. The spectrum (range 380-590 nm, resolution 1.5 nm) is very similar to a normal type-Ia supernova such as SNe 1981B or 1990N ∼ 10-14 days past maximum, based on the absence of the S II lines at 530.0 nm (seen in normal pre-maximum spectra) and the strengths of the Mg II, Fe II, and Na I features. The red colors previously reported for the supernova are consistent with a type-Ia supernova two weeks after maximum, and are not necessarily indicative of high reddening.”

D. Allen, Anglo-Australian Observatory; M. Burton and A. Walsh, University of New South Wales; and G. Stringfellow, University of California at Berkeley, report on near-infrared magnitudes of SN 1993L derived from images taken on May 1.79 UT with the IRIS camera on the Anglo-Australian Telescope: “After spline-fitting the underlying radiation from IC 5270, the magnitudes were J = 15.2, H = 13.7, K_{n} = 13.8. K_{n} is a narrow-band K filter covering 2.0-2.3 μm. The colors suggest a type-Ia supernova ∼ 3 weeks after maximum light.”

SUPERNOVA 1993J IN NGC 3031 X. Zhou, Beijing Astronomical Observatory, reports that further cor-

rections are needed to the photometry published by Zheng et al. on IAUC 5777, such that the following values must be added to the U, B, V, R, and I magnitudes, respectively, to place the data properly in the Johnson system: –0.06, –0.05, +0.04, –0.12, –0.21.

1993 May 3 Daniel W. E. Green